A football player throws a 400 g football at 30 de
Answer #1

E = 100 J

1/2 mV^2 = 100

V = 22.36 m/s

Vh = horizontal velocity = 19.36 m/s

Vv = vertical velocity = 11.18 m/s

horizontal distance = V^2sin(2*30) / g = 44.18 m

i) initial speed = 22.36 m/s

ii) Vv = vertical velocity = 11.18 m/s

iii) vertical acceleration will be downward due to gravity = g = 9.8 m/s^2

iv) t = Vsin(30) / g = 1.14 sec

v) T = 2t = 2.28 sec

vi) Vh = horizontal velocity = 19.36 m/s

vii) horizontal acceleration = 0